# Configuration spaces of circles in the plane

### v1

@@ Abstract.
We consider the space of all configurations of finitely many (potentially nested) circles in the plane. We prove that this space is aspherical, and compute the fundamental group of each of its connected components. It turns out these fundamental groups are obtained as iterated semdirect products of braid groups, with the structure for each component dictated by a finite rooted tree. These groups can be viewed as “braided” versions of the automorphism groups of such trees. We also discuss connections to statistical mechanics, topological data analysis, and geometric group theory.
@@ 1. Introduction
Configuration spaces are fundamental objects of study, with rich topological and algebraic properties. The prototypical configuration space—of $n$ indistinguishable points in the plane, written $\operatorname{Conf}_{n}(*,\mathbb{R}^{2})$ here—has been studied since at least the 1890s [^1] Perhaps unsurprisingly, Gauss is credited with work on braids dating back to the 1820s [Epp13, Fri19]., when Hurwitz considered the motion of points in the complex plane in order to identify group actions on more general Riemann surfaces [Hur91]. Fixing start and end configurations and studying how trajectories wind around each other (Figure 5) leads naturally to the fundamental group of $\operatorname{Conf}_{n}(*,\mathbb{R}^{2})$, which was presented algebraically by Artin in terms of the braid group $B_{n}$ [Art25, Art47]. Studying configurations of points in more general settings—arbitrary manifolds [FN62a], graphs [Abr00], and general stratified spaces [Pet17]—leads directly to some of the deepest ideas in modern mathematics; see [Knu, Kal] for surveys. Additionally, these ideas are finding new importance in modern engineering, with motion planning and robotics as particularly significant benefactors [BG18, MK20, Kas].
@@
In this paper we consider a simple variation of $\operatorname{Conf}_{n}(*,\mathbb{R}^{2})$ that studies configurations of circles instead of points. The resulting space of $n$ (unlabeled and non-intersecting) circles in the plane, written $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ here, is already more complex than its prototype because it is topologically disconnected. Indeed, any path between configurations of circles must preserve their so called nesting structure, which we now explain. By the Jordan Curve Theorem, any circle divides the plane into two connected components: one bounded (the inside) and one unbounded (the outside). Given two disjoint circles $C_{1}$ and $C_{2}$, $C_{1}$ is either contained in the inside or the outside of $C_{2}$. In the former case, as in Figure 1 (A), we say that $C_{1}$ is nested inside $C_{2}$. In the latter case, as in Figure 1 (B), we say that $C_{1}$ and $C_{2}$ are un-nested. This endows every configuration of $n$ circles with a circle containment order [SW88], also called a nesting poset structure in [CCF+20], which can be encoded into a tree $T$ with one root and $n$ other vertices. Intuitively, a path in $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ describes a way of moving the circles around without meeting each other, and this cannot change the tree $T$ associated to the configuration. See Section 3 for more details.
[…]
@@
Figure 1. Two configurations in different components of $\operatorname{Conf}_{2}(S^{1},\mathbb{R}^{2})$.
Our main result in this paper, Theorem A, includes the first (to our knowledge) explicit calculation of the fundamental group of each connected component of $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$. In the statement of this theorem, $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ is the connected component of $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ corresponding to the tree $T$, and $\operatorname{BAut}(T)$ is the iterated semidirect product of braid groups that serves as a “braided” version of $\operatorname{Aut}(T)$; see Definitions 3.9 and 4.4 for more.
@@
Let $T$ be a tree with $n$ non-root vertices. Then $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ is aspherical, and its fundamental group is isomorphic to $\operatorname{BAut}(T)$, hence it is a $K(\operatorname{BAut}(T),1)$.
Here, recall that a topological space $X$ is aspherical if $\pi_{k}(X)=0$ for all $k\neq 1$, and in this case it is called a $K(G,1)$ for its fundamental group.
@@ 1.2. Connections: topological data analysis, deep learning and sensor networks
Another motivation for our paper comes from a desire to quantify the expressive power of methods arising in applied topology. In Topological Data Analysis (TDA), one encounters the Reeb graph and the persistence barcode as featurization tools for shape analysis. The Reeb graph [Kro50, Ree46, Sha06] is a graph that tracks the evolution of connected components ($\pi_{0}$) of a space when viewed along a suitably tame function $f:X\to\mathbb{R}$. The barcode [CZCG04]—originally known as the persistence diagram [ELZ02]—measures the evolution of homology with field coefficients ($H_{k}$) when viewed along a similarly nice function. When $k=0$, and under certain technical assumptions [DeS21], one can view the barcode as a “linearized” version of the Reeb graph. However, as observed in [CCF+20], both the Reeb graph and the barcode are insensitive to certain embedding information, as illustrated in Figures 3 and 4.
@@
It should be noted that a variant of both the Reeb graph and the nesting poset called “the tree of shapes” [BCM03, CM10, WL22] has been used by the computer vision community to track nesting relationships among objects in an image. Indeed, many image segmentation algorithms proceed by finding steep changes in contrast (viewed as $||\nabla I||$ for $I$ some pixel intensity function) to identify the boundaries of objects. Heuristically, well-known results from differential topology, such as the Pre-Image Theorem [GP10, §4], provide theoretical guarantees that these boundaries should specify a topological arrangement of circles. Continuing this line of thought and motivated by current paradigms of geometric and topological deep learning [BBL+17, BBCV, PBB+24], one might ask for equivariance properties in a neural net whose input is an image and whose output is a tree of shapes. Our work provides a calculation of the groups that one can deform an image by and retain the same segmentation.
Finally, we note that our computation of $\pi_{1}(\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2}))$ should be applicable to the so called “pursuit-evasion problem” for mobile sensor networks [DSG06] by following recent work by Gunnar Carlsson and co-authors [CF, CFM], which computes $\pi_{0}$ of the space sections of a submersion $f:Z\to Y$ when $Y=[0,1]$ or $Y=S^{1}$ in terms of $\pi_{0}$ and $\pi_{1}$ of the fiber. Since we can think of sensor networks as covering some (potentially disconnected) two-manifold with boundary in $\mathbb{R}^{2}$, the complement (where intruders are allowed) should retract onto a configuration of circles in the plane. The map to $Y$ then encodes how this configuration evolves in time and our computation of $\pi_{1}$ offers a potential classification result for mobile sensor networks along with their evasion paths.
@@ 1.3. Connections: geometric group theory
Finally, another motivator for the present paper comes from geometric group theory. There has been a lot of recent interest in “braiding” groups that are intrinsically “permutational” in some way. The classical example is the braid groups themselves, which can be viewed as a braided version of the symmetric groups. Visually, this amounts to taking a picture for an element of $S_{n}$, viewed as $n$ arrows going from a line of $n$ dots to another line of $n$ dots, and introducing a specification of which arrows (now strands) go in front and which go behind as they cross. Algebraically, this can be realized by taking a group presentation in which each generator has a defining relation declaring it to have order 2, and simply removing these relations. Other, more recent examples of “braided” versions of groups include a variety of braided Thompson groups [Bri07, Deh06, BBCS08, FK08, FK11, BFM+16, Wit19, Spa21], braided Houghton groups [Deg00, GLU22], braided groups of automorphisms of an infinite, locally finite, regular tree [AC22, Cum, SW23, SZ24], and braided versions of self-similar groups and Röver–Nekrashevych groups [SZ24]. There is also a robust program originated by Jones, connecting Thompson’s groups to braids, knots, and links, and unitary representation theory [Jon17].
@@
In this paper, we “braid” the group of automorphisms of an arbitrary (not necessarily regular) finite, rooted tree $T$. We construct a braided automorphism group $\operatorname{BAut}(T)$, and prove it is isomorphic to the fundamental group of the connected component of the space of unlabeled configurations of circles in the plane corresponding to $T$ in a certain way. We also prove these spaces are aspherical, and so the space corresponding to $T$ is a $K(\operatorname{BAut}(T),1)$; recall that a $K(G,1)$, also called an Eilenberg–MacLane space for a group $G$, is a connected aspherical space whose fundamental group is $G$. (For discrete groups like $\operatorname{BAut}(T)$ this implies the space is also a classifying space for the group, in the sense of homotopy theory.) Finally, we prove similar results in the “pure braided” case, dealing with labeled configurations.
@@
One possible direction for future study is generalizing our results to locally finite infinite trees. With a good understanding of the groups $\operatorname{BAut}(T)$ for finite $T$, taking a limit should produce groups of the form $\operatorname{BAut}(T)$ for locally finite infinite $T$. This would generalize the situation from [SZ24] to trees that are not necessarily regular, and would likely produce new groups of interest in the world of big mapping class groups, as in [AV20]. Tying in to Subsection 1.1, we speculate that it might also lead to interesting variations on representation stability, as in [AM24].
@@ Acknowledgments
Thanks are due to Yuliy Baryshnikov, Matt Kahle, Robbie Lyman, and Thomas Rot for helpful discussions. Special thanks goes out to Renee Hoekzema for several enlightening research conversations that occurred in Oxford and Amsterdam during the first author’s sabbatical. JC’s research has been partially supported by NSF grant CCF-1850052 and NASA Contract 80GRC020C0016. MZ is supported by Simons grant #635763.
@@ 2. Configuration spaces of points, circles, and disks
Given a topological space $X$, the labeled configuration space of $n$ points in $X$ is the space $\operatorname{LConf}_{n}(*,X)$, whose elements are $n$-tuples $(x_{1},\dots,x_{n})$ of pairwise distinct elements $x_{i}\in X$, viewed as a subspace of $X^{n}$. There is a natural action of the symmetric group $S_{n}$ on $\operatorname{LConf}_{n}(*,X)$, given by permuting the entries in the tuples, and the orbit space $\operatorname{Conf}_{n}(*,X)$ is called the configuration space of $n$ points in $X$.
We are particularly interested in the case when $X=\mathbb{R}^{2}$. Recall that the fundamental group of $\operatorname{Conf}_{n}(*,\mathbb{R}^{2})$ is the $n$-strand braid group $B_{n}$ [FN62b], given by the presentation
@@
Visually, an element of $B_{n}=\pi_{1}(\operatorname{Conf}_{n}(*,\mathbb{R}^{2}))$ can be pictured by working in $\mathbb{R}^{2}\times[0,1]$, and drawing a “braid” consisting of a configuration of $n$ points for each moment of “time” in $[0,1]$. See Figure 5 for an example.
Figure 5. An element of $\pi_{1}(\operatorname{Conf}_{4}(*,\mathbb{R}^{2}))$, viewed as a braid.
We will also be concerned with configuration spaces of circles in the plane, and also of disks in the plane. By a circle we will always mean a subspace of $\mathbb{R}^{2}$, with the usual Euclidean metric, that is a circle of some positive radius centered at some point, and a disk will mean a closed disk with the same specifications. Configuration spaces of circles in $\mathbb{R}^{3}$ have been studied by Brendle–Hatcher in [BH13], see also [BWC07, Dam17], and are related to the so called loop braid groups. Here we look at circles in $\mathbb{R}^{2}$, where the most evident difference with the $\mathbb{R}^{3}$ case is that, as we will see, the space is disconnected.
@@
Let $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$ be the topological space whose elements are ordered $n$-tuples $(C_{1},\dots,C_{n})$ of pairwise disjoint circles $C_{i}$ in the plane. The topology on $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$ comes from realizing that each $C_{i}$ is specified by three data points, namely the coordinates of its center together with its radius, so we view $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$ as a subspace of $\mathbb{R}^{3n}$. We call $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$ the labeled configuration space of $n$ circles in the plane, where we view the circles as labeled by the numbers $1$ through $n$. There is a natural action of the symmetric group $S_{n}$ on $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$, given by permuting the entries in the tuples, and the orbit space $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ is called the (unlabeled) configuration space of $n$ circles in the plane. We can view elements of $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ as (unordered) sets $\{C_{1},\dots,C_{n}\}$ of $n$ pairwise disjoint circles in the plane. Call $\{C_{1},\dots,C_{n}\}$ the underlying set of $(C_{1},\dots,C_{n})$.
@@
Let $\operatorname{LConf}_{n}(D^{2},\mathbb{R}^{2})$ be the topological space whose elements are ordered $n$-tuples $(D_{1},\dots,D_{n})$ of pairwise disjoint disks $D_{i}$ in the plane. This can be viewed as a subspace of $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$, via the embedding that ignores the interiors of all the disks (yielding the subspace in which none of the circles are nested in each other). This is the labeled configuration space of $n$ disks in the plane, and we similarly define the (unlabeled) configuration space of $n$ disks in the plane to be the orbit space under the $S_{n}$-action.
As mentioned, the topology on $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$ and $\operatorname{LConf}_{n}(D^{2},\mathbb{R}^{2})$ comes from viewing them as subspaces of $\mathbb{R}^{3n}$. In fact, we can explicitly describe these subspaces, and see that they are semialgebraic sets, that is, they are defined as finite unions of sets of solutions to polynomial equalities and inequalities. Write the coordinates for elements of $\mathbb{R}^{3n}$ as $(r_{1},\dots,r_{n},x_{1},\dots,x_{n},y_{1},\dots,y_{n})$, with $r_{i}$ representing the radius of the $i$th circle and $(x_{i},y_{i})$ representing the center of the $i$th circle.
@@
The space $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$ is described by the semialgebraic set in $\mathbb{R}^{3n}$ consisting of all $3n$-tuples as above satisfying that $r_{i}>0$ for all $1\leq i\leq n$, and for all $1\leq i\neq j\leq n$ we have:
$(r_{i}-r_{j})^{2}-((x_{i}-x_{j})^{2}+(y_{i}-y_{j})^{2})>0\quad\text{ or }\quad(r_{i}+r_{j})^{2}-((x_{i}-x_{j})^{2}+(y_{i}-y_{j})^{2})<0\text{.}$
Moreover, $\operatorname{LConf}_{n}(D^{2},\mathbb{R}^{2})$ is the subspace where we insist on the second possibility for all $1\leq i\neq j\leq n$.
@@
The main goal of this paper is to understand the fundamental groups of these configuration spaces. The disk case turns out be essentially the same as the point case, and so yields the classical braid and pure braid groups, but the circle case is much more subtle.

### v2

@@ Abstract.
We consider the space of all configurations of finitely many (potentially nested) circles in the plane. We prove that this space is aspherical, and compute the fundamental group of each of its connected components. It turns out these fundamental groups are obtained as iterated semidirect products of subgroups of braid groups, with the structure for each component dictated by a finite rooted tree. These groups can be viewed as “braided” versions of the automorphism groups of such trees. We also discuss connections to statistical mechanics, topological data analysis, and geometric group theory.
@@ 1. Introduction
Configuration spaces are fundamental objects of study, with rich topological and algebraic properties. The prototypical configuration space—of $n$ indistinguishable points in the plane, written $\operatorname{UConf}_{n}(*,\mathbb{R}^{2})$ here—has been studied since at least the 1890s [^1] Perhaps unsurprisingly, Gauss is credited with work on braids dating back to the 1820s [Epp13, Fri19]., when Hurwitz considered the motion of points in the complex plane in order to identify group actions on more general Riemann surfaces [Hur91]. Fixing start and end configurations and studying how trajectories wind around each other (Figure 5) leads naturally to the fundamental group of $\operatorname{UConf}_{n}(*,\mathbb{R}^{2})$, which was presented algebraically by Artin in terms of the braid group $B_{n}$ [Art25, Art47]. Studying configurations of points in more general settings—arbitrary manifolds [FN62a], graphs [Abr00], and general stratified spaces [Pet17]—leads directly to some of the deepest ideas in modern mathematics; see [Knu, Kal25] for surveys. Additionally, these ideas are finding new importance in modern engineering, with motion planning and robotics as particularly significant benefactors [BG18, MK20, Kas].
@@
In this paper we consider a simple variation of $\operatorname{UConf}_{n}(*,\mathbb{R}^{2})$ that studies configurations of circles instead of points. The resulting space of $n$ unlabeled and non-intersecting circles in the plane, written $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ here, is already more complex than its prototype because it is topologically disconnected. Indeed, any path between configurations of circles must preserve their so called nesting structure, which we now explain. By the Jordan Curve Theorem, any circle divides the plane into two connected components: one bounded (the inside) and one unbounded (the outside). Given two disjoint circles $C_{1}$ and $C_{2}$, $C_{1}$ is either contained in the inside or the outside of $C_{2}$. In the former case, as in Figure 1 (A), we say that $C_{1}$ is nested inside $C_{2}$. If neither circle is nested in the other, as in Figure 1 (B), we say that $C_{1}$ and $C_{2}$ are un-nested. This endows every configuration of $n$ circles with a circle containment order [SW88], also called a nesting poset structure in [CCF+20], which can be encoded in a tree $T$ with one root and $n$ other vertices—an observation well understood by students of the little disks operad, cf. [Idr22, §5.3.1]. As a brief aside, we note that beyond operads, trees also arise naturally in the Fulton–MacPherson compactification of $\operatorname{UConf}_{n}(*,\mathbb{R}^{2})$; see [FM94], and also [Sin04] for a more modern perspective. We expect that the simultaneous appearance of these trees is not a coincidence, but leave the precise connection for future research. For us, the tree $T$ simply encodes the containment order of circles, and a path in $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ cannot change this containment order, thus making the tree $T$ an invariant of a connected component. See Section 3 for more details. As a final remark, moving up a dimension, the space $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{3})$ of configurations of $n$ circles in $3$-space is a well-studied object, related to so called loop braid groups, see for example [Gol81, BWC07, BH13, Dam17]. Also see [BB25, BB, DK19] for more in this direction.
$C_{1}$$C_{2}$
@@
Figure 1. Two configurations in different components of $\operatorname{UConf}_{2}(S^{1},\mathbb{R}^{2})$.
Our main result in this paper, Theorem A, includes the first (to our knowledge) explicit calculation of the fundamental group of each connected component of $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$. In the statement of this theorem, $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$ is the connected component of $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ corresponding to the tree $T$, and $\operatorname{bAut}(T)$ is the iterated semidirect product of braid groups that serves as a “braided” version of $\operatorname{Aut}(T)$; see Definitions 3.9 and 4.4 for more.
@@
Let $T$ be a tree with $n$ non-root vertices. Then $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$ is aspherical, and its fundamental group is isomorphic to $\operatorname{bAut}(T)$, hence it is a $K(\operatorname{bAut}(T),1)$.
Here, recall that a topological space $X$ is aspherical if $\pi_{k}(X)=0$ for all $k\neq 1$, and in this case it is called a $K(G,1)$ for its fundamental group $G$.
@@ 1.2. Connections: topological data analysis, deep learning and sensor networks
Another motivation for our paper comes from a desire to quantify the expressive power of methods arising in applied topology. In Topological Data Analysis (TDA), one encounters the Reeb graph and the persistence barcode as featurization tools for shape analysis. The Reeb graph [Kro50, Ree46, Sha06] is a graph that tracks the evolution of connected components ($\pi_{0}$) of a space when viewed along a suitably tame function $f\colon X\to\mathbb{R}$. The barcode [CZCG04]—originally known as the persistence diagram [ELZ02]—measures the evolution of homology with field coefficients ($H_{k}$) when viewed along a similarly nice function. When $k=0$, and under certain technical assumptions [DeS21], one can view the barcode as a “linearized” version of the Reeb graph. However, as observed in [CCF+20], both the Reeb graph and the barcode are insensitive to certain embedding information, as illustrated in Figures 3 and 4.
@@
It should be noted that a variant of both the Reeb graph and the nesting poset called “the tree of shapes” [BCM03, CM10, WL22] has been used by the computer vision community to track nesting relationships among objects in an image. Indeed, many image segmentation algorithms proceed by finding steep changes in contrast (viewed as $\nabla I||$ for $I$ some pixel intensity function) to identify the boundaries of objects. Heuristically, well-known results from differential topology, such as the Pre-Image Theorem [GP10, §4], provide theoretical guarantees that these boundaries should specify a topological arrangement of circles. Continuing this line of thought and motivated by current paradigms of geometric and topological deep learning [BBL+17, BBCV, PBB+24], one might ask for equivariance properties in a neural net whose input is an image and whose output is a tree of shapes. Our work provides a calculation of the groups that one can deform an image by and retain the same segmentation.
Finally, we note that our computation of $\pi_{1}(\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2}))$ should be applicable to the so called “pursuit-evasion problem” for mobile sensor networks [DSG06] by following recent work by Gunnar Carlsson and co-authors [CF, CFM], which computes $\pi_{0}$ of the space sections of a submersion $f\colon Z\to Y$ when $Y=[0,1]$ or $Y=S^{1}$ in terms of $\pi_{0}$ and $\pi_{1}$ of the fiber. Since we can think of sensor networks as covering some (potentially disconnected) two-manifold with boundary in $\mathbb{R}^{2}$, the complement (where intruders are allowed) should retract onto a configuration of circles in the plane. The map to $Y$ then encodes how this configuration evolves in time and our computation of $\pi_{1}$ offers a potential classification result for mobile sensor networks along with their evasion paths.
@@ 1.3. Connections: geometric group theory
Finally, another motivator for the present paper comes from geometric group theory. There has been a lot of recent interest in “braiding” groups that are intrinsically “permutational” in some way. The classical example is the braid groups themselves, which can be viewed as a braided version of the symmetric groups. Visually, this amounts to taking a picture for an element of $S_{n}$, viewed as $n$ arrows going from a line of $n$ dots to another line of $n$ dots, and introducing a specification of which arrows (now strands) go in front and which go behind as they cross. Algebraically, this can be realized by taking a group presentation in which each generator has a defining relation declaring it to have order 2, and simply removing these relations. Other, more recent examples of “braided” versions of groups include a variety of braided Thompson groups [Bri07, Deh06, BBCS08, FK08, FK11, BFM+16, Wit19, Spa21], braided Houghton groups [Deg00, GLU22], braided groups of automorphisms of an infinite, locally finite, regular tree [AC22, Cum24, SW23, SZ24], and braided versions of self-similar groups and Röver–Nekrashevych groups [SZ24]. There is also a robust program originated by Jones, connecting Thompson’s groups to braids, knots, and links, and unitary representation theory [Jon17].
@@
In this paper, we “braid” the group of automorphisms of an arbitrary (not necessarily regular) finite, rooted tree $T$. We construct a braided automorphism group $\operatorname{bAut}(T)$, and prove it is isomorphic to the fundamental group of the connected component of the space of unlabeled configurations of circles in the plane corresponding to $T$ in a certain way. We also prove these spaces are aspherical, and so the space corresponding to $T$ is a $K(\operatorname{bAut}(T),1)$. (For discrete groups like $\operatorname{bAut}(T)$ this implies the space is also a classifying space for the group, in the sense of homotopy theory.) Finally, we prove similar results in the “pure braided” case, dealing with labeled configurations. As a remark, the group $\operatorname{bAut}(T)$ has appeared before in a different guise in work of Kordek and Margalit [KM22], as the group of mapping classes stabilizing a multicurve in a disk with some number of marked points. Here the tree $T$ comes from the nesting structure of the multicurve. It would be interesting to confirm that this is not a coincidence, for example using a generalization of the Birman exact sequence in Theorem 9.1 of [FM12].
@@
Another possible direction for future study involves generalizing our results to locally finite infinite trees. With a good understanding of the groups $\operatorname{bAut}(T)$ for finite $T$, taking a limit should produce groups of the form $\operatorname{bAut}(T)$ for locally finite infinite $T$. This would generalize the situation from [SZ24] to trees that are not necessarily regular, and would likely produce new groups of interest in the world of big mapping class groups, as in [AV20]. Tying in to Subsection 1.1, we speculate that it might also lead to interesting variations on representation stability, as in [AM24]. One could also consider configurations of circles on surfaces other than $\mathbb{R}^{2}$, such as the $2$-sphere, torus, surface of genus $g$ for $g>1$, and so forth.
@@ Acknowledgments
Thanks are due to Yuliy Baryshnikov, Matt Kahle, Robbie Lyman, Dan Margalit, and Thomas Rot for helpful discussions. Special thanks goes out to Renee Hoekzema for several enlightening research conversations that occurred in Oxford and Amsterdam during the first author’s sabbatical. We are also grateful to the referee for several very helpful suggestions. JC’s research has been partially supported by NSF grant CCF-1850052 and NASA Contract 80GRC020C0016. MZ is supported by Simons grant #635763.
@@ 2. Configuration spaces of points, circles, and disks
Given a topological space $X$, the (labeled) configuration space of $n$ points in $X$ is the space $\operatorname{Conf}_{n}(*,X)$, whose elements are $n$-tuples $(x_{1},\dots,x_{n})$ of pairwise distinct elements $x_{i}\in X$, viewed as a subspace of $X^{n}$. There is a natural action of the symmetric group $S_{n}$ on $\operatorname{Conf}_{n}(*,X)$, given by permuting the entries in the tuples, and the orbit space $\operatorname{UConf}_{n}(*,X)$ is called the unlabeled configuration space of $n$ points in $X$.
We are particularly interested in the case when $X=\mathbb{R}^{2}$. Recall that the fundamental group of $\operatorname{UConf}_{n}(*,\mathbb{R}^{2})$ is the $n$-strand braid group $B_{n}$ [FN62b], given by the presentation
@@
Visually, an element of $B_{n}=\pi_{1}(\operatorname{UConf}_{n}(*,\mathbb{R}^{2}))$ can be pictured by working in $\mathbb{R}^{2}\times[0,1]$, and drawing a “braid” consisting of a configuration of $n$ points for each moment of “time” in $[0,1]$. See Figure 5 for an example.
Figure 5. An element of $\pi_{1}(\operatorname{UConf}_{4}(*,\mathbb{R}^{2}))$, viewed as a braid.
We will also be concerned with configuration spaces of circles in the plane, and also of disks in the plane. By a circle we will always mean a subspace of $\mathbb{R}^{2}$, with the usual Euclidean metric, that is a circle of some positive radius centered at some point, and a disk will mean a closed disk with the same specifications. Configuration spaces of circles in $\mathbb{R}^{3}$ have been studied by Brendle–Hatcher in [BH13], see also [BWC07, Dam17, Gol81], and are related to the so called loop braid groups. Here we look at circles in $\mathbb{R}^{2}$, where the most evident difference with the $\mathbb{R}^{3}$ case is that, as we will see, the space is disconnected.
@@
Let $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ be the topological space whose elements are ordered $n$-tuples $(C_{1},\dots,C_{n})$ of pairwise disjoint circles $C_{i}$ in the plane. The topology on $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ comes from realizing that each $C_{i}$ is specified by three data points, namely the coordinates of its center together with its radius, so we view $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ as a subspace of $\mathbb{R}^{3n}$. We call $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ the (labeled) configuration space of $n$ circles in the plane, where we view the circles as labeled by the numbers $1$ through $n$. There is a natural action of the symmetric group $S_{n}$ on $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$, given by permuting the entries in the tuples, and the orbit space $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ is called the unlabeled configuration space of $n$ circles in the plane. We can view elements of $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ as (unordered) sets $\{C_{1},\dots,C_{n}\}$ of $n$ pairwise disjoint circles in the plane. Call $\{C_{1},\dots,C_{n}\}$ the underlying set of $(C_{1},\dots,C_{n})$.
@@
Let $\operatorname{Conf}_{n}(D^{2},\mathbb{R}^{2})$ be the topological space whose elements are ordered $n$-tuples $(D_{1},\dots,D_{n})$ of pairwise disjoint disks $D_{i}$ in the plane. This can be viewed as a subspace of $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$, via the embedding that ignores the interiors of all the disks (yielding the subspace in which none of the circles are nested in each other). This is the (labeled) configuration space of $n$ disks in the plane, and we similarly define the unlabeled configuration space of $n$ disks in the plane $\operatorname{UConf}_{n}(D^{2},\mathbb{R}^{2})$ to be the orbit space under the $S_{n}$-action.
As mentioned, the topology on $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ and $\operatorname{Conf}_{n}(D^{2},\mathbb{R}^{2})$ comes from viewing them as subspaces of $\mathbb{R}^{3n}$. In fact, we can explicitly describe these subspaces, and see that they are semialgebraic sets, that is, they are defined as finite unions of sets of solutions to polynomial equalities and inequalities. Write the coordinates for elements of $\mathbb{R}^{3n}$ as $(r_{1},\dots,r_{n},x_{1},\dots,x_{n},y_{1},\dots,y_{n})$, with $r_{i}$ representing the radius of the $i$th circle and $(x_{i},y_{i})$ representing the center of the $i$th circle.
@@
The space $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ is described by the semialgebraic set in $\mathbb{R}^{3n}$ consisting of all $3n$-tuples as above satisfying that $r_{i}>0$ for all $1\leq i\leq n$, and for all $1\leq i\neq j\leq n$ we have:
[…]
Moreover, $\operatorname{Conf}_{n}(D^{2},\mathbb{R}^{2})$ is the subspace where we insist on the second possibility for all $1\leq i\neq j\leq n$.
@@
The main goal of this paper is to understand the fundamental groups of these labeled and unlabeled configuration spaces. The disk case turns out to be essentially the same as the point case, and so yields the classical braid and pure braid groups, but the circle case is much more subtle.